# Separating Tongue from Thought: Activation Patching Reveals Language-Agnostic Concept Representations in Transformers

### v3

@@ 6 Generating descriptions for latents
We just showed that LLMs can effectively translate concept representations averaged across multiple languages, providing evidence for language-agnostic concept representations. In this section, we explore whether these mean representations, which theoretically capture language-independent concepts, can be described by the model in natural language as effectively as concepts expressed in a single language.
@@ 6.1 Definition prompt
where $C^{\ell}_{1},\dots,C^{\ell}_{n}$ are concepts in language $\ell$ and $D^{\ell}_{1},\dots,D^{\ell}_{n}$ are their descriptions in language $\ell$. We denote $\text{DP}^{\text{collect}}(\ell,C)$ the prompt template that ends at the last token of $C^{\ell}$.
@@ 6.2 Patching setup
    $\mathcal{S}_{\text{trans}}=\{\text{TP}^{\text{collect}}(\ell_{(\text{in})}^{i},\ell_{S}^{i},C_{S})\}_{i\in\{1,\dots,n\}}.$
@@
    $\mathcal{S}_{\text{def}}=\{\text{DP}^{\text{collect}}(\ell_{S}^{i},C_{S})\}_{i\in\{1,\dots,n\}}.$
@@ 6.3 Experiment
To evaluate the different definitions, we use the sentence-transformers library Reimers and Gurevych (2019) to run paraphrase-multilingual-mpnet-base-v2555https://huggingface.co/sentence-transformers/paraphrase-multilingual-mpnet-base-v2, a semantic similarity model distilled from Song et al. (2020) using the methods from Reimers and Gurevych (2020). To assess the quality of the generated definition, we compute the similarity score between embeddings of the definition and the mean embedding of the ground truth definitions [^6] To be able to compare the score of the generated definition with the ground truth baseline, we compute its similarity score with the mean embedding of the other ground truth definitions - excluding the one that was randomly chosen as the ground truth baseline.
@@ 6.4 Results
Other models. In Appendix C.3 we show that those findings generalize to other models and languages.
@@ 7 Conclusion
In this paper, we showed that transformers use language-agnostic latent representations of concepts when processing word-level translation prompts. We achieved this by patching latents between parallel forward passes for translation prompts that differed in both input and output languages, as well as in the specific concepts being translated. Our main finding was that translation performance improves when the transformer is forced to translate a concept representation averaged across multiple languages. This finding speaks for language-agnostic concept representations. As we argued, for language-agnostic concept representations, taking the mean representation of a concept across languages should not impair the LLM’s ability to translate this concept. In contrast, for language-specific ones, taking the mean should result in interference between the different language-specific versions of the concept. Thus, our results are consistent with findings from previous work Wendler et al. (2024) indicating that Llama-2 represents concepts in a concept space independent of the language of the prompt. Our work also provides evidence that findings from BERT models Wu et al. (2019); Pires et al. (2019) generalize to a wide range of decoder-only transformers.
@@ Appendix A Patchscope experiment
Appendix B Random prompt task experiment
In order to investigate the leftmost part of Figure 7(a) more deeply, we performed additional experiments in which we explore “random" source prompts instead of translation source prompts.
The experimental setting here is similar to the one in Sec. 4, except for the fact that instead of patching in latents from a translation source prompt we patch latents from different “random" source prompts. For the random source prompts, we gradually move away from the prompting template.
@@ Appendix B Random prompt task experiment
(c) Random prompt with “@" instead of quotation mark
@@
Figure 8: (a) activation patching experiment with a randomized source prompt (random concepts, and languages, but same template) and a fr to zh target prompt. (b) we construct a source prompt with empty context. (c) we replace the quotation mark with @ in the random source prompt from (a). (d) we randomly shuffle the source prompts from (c). We patch at the last token respectively. For each of the plots, the x-axis shows at which layer the patching was performed during the forward pass on the target prompt and the y-axis shows the probability of predicting the correct concept in language $\ell$ (see legend). We only plot the target (“tgt") concept, as there is no source concept to predict. We report means and 95% Gaussian confidence intervals computed over 200 source-, target prompt pairs.
Same template. In Figure 8(a), we randomized both input and output language as well as concepts in the source prompts, resulting in prompts of the following form: A: “$\text{{{cat}}}^{\text{{{de}}}}$" - B: “$\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}{{dog}}}}^{\text{{{it}}}}$" A: “$\text{{{owl}}}^{\text{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}{{ja}}}}$" - B: “$\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}{{sun}}}}^{\text{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}{{hi}}}}$" A: “$\text{{{ice}}}^{\text{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}{{fr}}}}$" - B: “
@@
Empty context. For example, replacing the source prompt with an empty prompt, merely containing B: " results in Figure 8(b). In contrast to Figure 8(a), the target concept in target language probability drops already starting from layer 4. We think this is due to the fact that until layer 4 the quotation mark token information which is shared among the two prompting templates “dominates” the latent representation and is not yet converted to a task specific position marker yet. Then, starting from layer 4 the latent representation of the last token also aggregates task specific information, in particular, the fact that the quotation mark in this task actually marks the position after which the translated word should be decoded. As a result, replacing the task specific quotation mark embedding, which contains the information that the translated word comes next, with the “empty-context”-one, which does not contain this information, results in a performance drop.
Modified template. Next, replacing the quotation marks by “@" (Figure 8(c)) in the random prompt, i.e., A: @$\text{{{cat}}}^{\text{{{de}}}}$@ - B: @$\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}{{dog}}}}^{\text{{{it}}}}$@ A: @$\text{{{owl}}}^{\text{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}{{ja}}}}$@ - B: @$\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}{{sun}}}}^{\text{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}{{hi}}}}$@ A: @$\text{{{ice}}}^{\text{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}{{fr}}}}$@ - B: @
@@
Appendix C Other models and languages
In this section, we report results for additional models, namely, Mistral 7B Jiang et al. (2023), Llama 3 8B Dubey et al. (2024), Qwen 1.5 7B Bai et al. (2023) and Llama 2 70B Touvron et al. (2023). We also include Aya 23 8B Aryabumi et al. (2024) for the mean patching experiment in App C.2.
C.1 Exploratory analysis
@@
Figure 9: Our first patching experiment with a de to it source prompt and a fr to zh target prompt with different concepts. We patch at the last token. For each of the plots the x-axis shows at which layer the patching was performed during the forward pass on the target prompt and the y-axis shows the probability of predicting the correct concept in language $\ell$ (see legend). In the legend the prefix “src" stands for source and “tgt" for target concept. The orange dashed line and blue dash-dotted line correspond to the mean accuracy on source and target prompt. We report means and 95% Gaussian confidence intervals computed over 200 source-, target prompt pairs featuring 41 source concepts and 38 target concepts.
C.2 Ruling out hypotheses
@@
In addition, instead of just patching in the mean over different language pairs (Figure 11c, 12c), we also patch in the mean over contexts composed of different concept words in Figure 11b, 12b. In particular, we take the mean over 5 different few-shot contexts from the same language pair. E.g.: Deutsch: “Dorf" - Italiano: “villaggio" Deutsch: “Buch
Deutsch: “Zitrone" - Italiano: “limone" Deutsch: “Buch
@@
We also provide some extra languages in Figure 10
Figure 10: multiple languages with 200 pairs each and 95% confidence interval.
@@
Figure 11: Here we use different input languages (de, fr), different concepts, different output languages (it, zh) in (a). In (b) we use the same source and target language pairs as in (a). In (c) we use multiple source input languages de, nl, zh, es, ru and output languages it, fi, es, ru, ko. We patch at the last token of the concept-word at all layers from $j$ to $31$. In (a) we patch latents from the single source prompt. In (b) for each concept, we patch the average latent over different few-shot de to it translation contexts. In (c) we patch the mean of the latents over the source prompts. For each of the plots, the x-axis shows at which layer the patching was performed during the forward pass on the target prompt and the y-axis shows the probability of predicting the correct concept in language $\ell$ (see legend). The prefix “src" stands for source and “tgt" for target concept. We report means and 95% Gaussian confidence intervals computed over a dataset of size 200.
@@
Gemma-2 2B
Figure 12: Here we use different input languages (de, fr), different concepts, different output languages (it, zh) in (a). In (b) we use the same source and target language pairs as in (a). In (c) we use multiple source input languages de, nl, zh, es, ru and output languages it, fi, es, ru, ko. We patch at the last token of the concept-word at all layers from $j$ to $31$. In (a) we patch latents from the single source prompt. In (b) for each concept, we patch the average latent over different few-shot de to it translation contexts. In (c) we patch the mean of the latents over the source prompts. For each of the plots, the x-axis shows at which layer the patching was performed during the forward pass on the target prompt and the y-axis shows the probability of predicting the correct concept in language $\ell$ (see legend). The prefix “src" stands for source and “tgt" for target concept. We report means and 95% Gaussian confidence intervals computed over a dataset of size 200.
C.3 Similarity comparison

### v4

@@ 6 Generating descriptions for latents
We just showed that LLMs can effectively translate concept representations averaged across multiple languages, providing evidence for language-agnostic concept representations. In this section, we explore whether these mean representations, which theoretically capture language-independent concepts, can be described by the model in natural language as effectively as concepts expressed in a single language. This approach builds on the definition modeling paradigm Noraset et al. (2017); Mickus et al. (2022); Chang and Chen (2019); Gardner et al. (2022), which uses natural language generation as a transparent evaluation of semantic representations.
@@ 6.1 Definition prompt
where $C^{\ell}_{1},\dots,C^{\ell}_{n}$ are concepts in language $\ell$ and $D^{\ell}_{1},\dots,D^{\ell}_{n}$ are their descriptions in language $\ell$. We denote $\text{DP}^{\text{concept}}(\ell,C)$ the prompt template that ends at the last token of $C^{\ell}$. For example, $\text{DP}^{\text{concept}}(\textsc{{en}},\textsc{{ant}})$ could be:
“apple” : “Fruit with red or yellow or green skin” “ant
@@ 6.2 Patching setup
    $\mathcal{S}_{\text{trans}}=\{\text{TP}^{\text{concept}}(\ell_{(\text{in})}^{i},\ell_{S}^{i},C_{S})\}_{i\in\{1,\dots,n\}}.$
@@
    $\mathcal{S}_{\text{def}}=\{\text{DP}^{\text{concept}}(\ell_{S}^{i},C_{S})\}_{i\in\{1,\dots,n\}}.$
@@ 6.3 Experiment
To evaluate the different definitions, we use the sentence-transformers library Reimers and Gurevych (2019) to run paraphrase-multilingual-mpnet-base-v2999https://huggingface.co/sentence-transformers/paraphrase-multilingual-mpnet-base-v2, a semantic similarity model distilled from Song et al. (2020) using the methods from Reimers and Gurevych (2020). To assess the quality of the generated definition, we compute the similarity score between embeddings of the definition and the mean embedding of the ground truth definitions [^10] To be able to compare the score of the generated definition with the ground truth baseline, we compute its similarity score with the mean embedding of the other ground truth definitions - excluding the one that was randomly chosen as the ground truth baseline.
@@ 6.4 Results
Other models. In Appendix D.3 we show that those findings generalize to other models and languages.
@@ 7 Conclusion
In this paper, we showed that transformers use language-agnostic latent representations of concepts when processing word-level translation prompts. We achieved this by patching latents between parallel forward passes for translation prompts that differed in both input and output languages, as well as in the specific concepts being translated. Our main finding was that translation performance improves when the transformer is forced to translate a concept representation averaged across multiple languages. This finding speaks for language-agnostic concept representations. As we argued, for language-agnostic concept representations, taking the mean representation of a concept across languages should not impair the LLM’s ability to translate this concept. In contrast, for language-specific ones, taking the mean should result in interference between the different language-specific versions of the concept. Thus, our results are consistent with findings from previous work Wendler et al. (2024) indicating that Llama-2 represents concepts in a concept space independent of the language of the prompt. Our work also provides evidence that findings from BERT models Conneau et al. (2020b); Pires et al. (2019) generalize to a wide range of decoder-only transformers. Our findings open several important avenues for future research. Understanding these disentangled representations could improve cross-lingual transfer learning and reduce the computational costs of multilingual training by leveraging shared concept spaces more efficiently. Moreover, our results provide mechanistic insights into the root causes of Western cultural biases in multilingual LLMs, suggesting that biases may propagate through the shared concept space. We hope these contributions will guide efforts to build more controllable, efficient, and culturally-aware multilingual language models.
@@ Appendix A Patchscope experiment
Appendix B Translation Pair Construction
To ensure reproducibility of our experiments, we provide the pseudocode for constructing translation pairs used in our activation patching experiments. The complete implementation is available in our codebase at notebooks/obj_patch_translation.ipynb.
Algorithm 1 Construction of Translation Pairs for Patching Experiments
1:Set of source languages $\mathcal{L}_{S}$, target language $\ell_{T}$, number of pairs $n$
2:Set of valid translation pairs $\mathcal{P}$
3:Load BabelNet translations for all languages
4:$\mathcal{T}_{S}\leftarrow$ GetTranslations($\mathcal{L}_{S}$)
5:$\mathcal{T}_{T}\leftarrow$ GetTranslations($\ell_{T}$)
6:$\mathcal{P}\leftarrow\emptyset$triangleright Initialize valid pairs
7:$\mathcal{A}\leftarrow$ Combinations($\mathcal{T}_{S},\mathcal{T}_{T}$) $\triangleright$ All possible pairs
8:Shuffle($\mathcal{A}$) $\triangleright$ Randomize order
9:for $(w_{S},w_{T})\in\mathcal{A}$ do
10:     if $\text{concept}(w_{S})=\text{concept}(w_{T})$ then
11:         continue $\triangleright$ Skip same concepts
12:     end if
13:     if HasTokenCollisions($w_{S},w_{T}$then
14:         continue $\triangleright$ Skip pairs with token overlap
15:     end if
16:     $\mathcal{P}\leftarrow\mathcal{P}\cup\{(w_{S},w_{T})\}$
17:     if $|\mathcal{P}|\geq n$ then
18:         break $\triangleright$ Sufficient pairs collected
19:     end if
20:end for
21:return $\mathcal{P}$
Key constraints:
  • Different concepts: We ensure that source and target words represent different concepts to enable meaningful patching experiments.
  • No token collisions: As described in Section 3, we track sets of tokens $w(C^{\ell})$ for each concept-language pair. To ensure clean probability measurements, we verify that there is no overlap between the token sets of paired concepts across all languages used in the experiment.
  • Randomization: Pairs are shuffled before selection to avoid systematic biases in concept or language selection.
For our experiments, we typically use $n=200$ pairs, with concepts drawn from the 200 picturable words from the Basic English word list, ensuring sufficient statistical power while maintaining computational feasibility.
Appendix C Random prompt task experiment
In order to investigate the leftmost part of Figure 7(a) more deeply, we performed additional experiments in which we explore “random” source prompts instead of translation source prompts.
The experimental setting here is similar to the one in Sec. 4, except for the fact that instead of patching in latents from a translation source prompt we patch latents from different “random” source prompts. For the random source prompts, we gradually move away from the prompting template.
@@ Appendix C Random prompt task experiment
(c) Random prompt with “@” instead of quotation mark
@@
Figure 8: (a) activation patching experiment with a randomized source prompt (random concepts, and languages, but same template) and a fr to zh target prompt. (b) we construct a source prompt with empty context. (c) we replace the quotation mark with @ in the random source prompt from (a). (d) we randomly shuffle the source prompts from (c). We patch at the last token respectively. For each of the plots, the x-axis shows at which layer the patching was performed during the forward pass on the target prompt and the y-axis shows the probability of predicting the correct concept in language $\ell$ (see legend). We only plot the target (“tgt”) concept, as there is no source concept to predict. We report means and 95% Gaussian confidence intervals computed over 200 source-, target prompt pairs.
Same template. In Figure 8(a), we randomized both input and output language as well as concepts in the source prompts, resulting in prompts of the following form: A: “$\text{{{cat}}}^{\text{{{de}}}}$” - B: “$\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}{{dog}}}}^{\text{{{it}}}}$ A: “$\text{{{owl}}}^{\text{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}{{ja}}}}$” - B: “$\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}{{sun}}}}^{\text{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}{{hi}}}}$ A: “$\text{{{ice}}}^{\text{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}{{fr}}}}$” - B: “
@@
Empty context. For example, replacing the source prompt with an empty prompt, merely containing B: “ results in Figure 8(b). In contrast to Figure 8(a), the target concept in target language probability drops already starting from layer 4. We think this is due to the fact that until layer 4 the quotation mark token information which is shared among the two prompting templates “dominates” the latent representation and is not yet converted to a task specific position marker yet. Then, starting from layer 4 the latent representation of the last token also aggregates task specific information, in particular, the fact that the quotation mark in this task actually marks the position after which the translated word should be decoded. As a result, replacing the task specific quotation mark embedding, which contains the information that the translated word comes next, with the “empty-context”-one, which does not contain this information, results in a performance drop.
Modified template. Next, replacing the quotation marks by “@” (Figure 8(c)) in the random prompt, i.e., A: @$\text{{{cat}}}^{\text{{{de}}}}$@ - B: @$\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}{{dog}}}}^{\text{{{it}}}}$@ A: @$\text{{{owl}}}^{\text{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}{{ja}}}}$@ - B: @$\text{{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}{{sun}}}}^{\text{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}{{hi}}}}$@ A: @$\text{{{ice}}}^{\text{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}{{fr}}}}$@ - B: @
@@
Appendix D Other models and languages
In this section, we report results for additional models, namely, Mistral 7B Jiang et al. (2023), Llama 3 8B Dubey et al. (2024), Qwen 1.5 7B Bai et al. (2023) and Llama 2 70B Touvron et al. (2023). We also include Aya 23 8B Aryabumi et al. (2024) for the mean patching experiment in App D.2.
D.1 Exploratory analysis
@@
Figure 9: Our first patching experiment with a de to it source prompt and a fr to zh target prompt with different concepts. We patch at the last token. For each of the plots the x-axis shows at which layer the patching was performed during the forward pass on the target prompt and the y-axis shows the probability of predicting the correct concept in language $\ell$ (see legend). In the legend the prefix “src” stands for source and “tgt” for target concept. The orange dashed line and blue dash-dotted line correspond to the mean accuracy on source and target prompt. We report means and 95% Gaussian confidence intervals computed over 200 source-, target prompt pairs featuring 41 source concepts and 38 target concepts.
D.2 Ruling out hypotheses
@@
In addition, instead of just patching in the mean over different language pairs (Figure 11c, 12c), we also patch in the mean over contexts composed of different concept words in Figure 11b, 12b. In particular, we take the mean over 5 different few-shot contexts from the same language pair. E.g.: Deutsch: “Dorf” - Italiano: “villaggio” Deutsch: “Buch
Deutsch: “Zitrone” - Italiano: “limone” Deutsch: “Buch
@@
Figure 10: Mean patching experiment replicated on multiple languages with 200 pairs each and 95% confidence interval.
@@
Figure 11: Here we use different input languages (de, fr), different concepts, different output languages (it, zh) in (a). In (b) we use the same source and target language pairs as in (a). In (c) we use multiple source input languages de, nl, zh, es, ru and output languages it, fi, es, ru, ko. We patch at the last token of the concept-word at all layers from $j$ to $31$. In (a) we patch latents from the single source prompt. In (b) for each concept, we patch the average latent over different few-shot de to it translation contexts. In (c) we patch the mean of the latents over the source prompts. For each of the plots, the x-axis shows at which layer the patching was performed during the forward pass on the target prompt and the y-axis shows the probability of predicting the correct concept in language $\ell$ (see legend). The prefix “src” stands for source and “tgt” for target concept. We report means and 95% Gaussian confidence intervals computed over a dataset of size 200.
@@
Gemma 2B
Figure 12: Here we use different input languages (de, fr), different concepts, different output languages (it, zh) in (a). In (b) we use the same source and target language pairs as in (a). In (c) we use multiple source input languages de, nl, zh, es, ru and output languages it, fi, es, ru, ko. We patch at the last token of the concept-word at all layers from $j$ to $31$. In (a) we patch latents from the single source prompt. In (b) for each concept, we patch the average latent over different few-shot de to it translation contexts. In (c) we patch the mean of the latents over the source prompts. For each of the plots, the x-axis shows at which layer the patching was performed during the forward pass on the target prompt and the y-axis shows the probability of predicting the correct concept in language $\ell$ (see legend). The prefix “src” stands for source and “tgt” for target concept. We report means and 95% Gaussian confidence intervals computed over a dataset of size 200.
D.3 Similarity comparison